OpenStudy (anonymous):

please verify the following using double-angle identities (no shortcuts please): \[\tan \theta = (\sin 2\theta)/(1 + \cos 2\theta)\] \[\tan \theta = (1 - \cos 2\theta)/(\sin 2\theta)\] \[(2\cos 2\theta)/(1 + \cos 2\theta) = 1 - \tan^{2} \theta\]

OpenStudy (anonymous):

left side only please :)

OpenStudy (anonymous):

left side is just the reverse of right side...if you can take the right side and make it into the left, then just reverse the work and you have it the other way... To help you out, however, i'll do one left side :)

OpenStudy (amistre64):

i try to put it all into sin and cos to see things better

OpenStudy (anonymous):

Wot, no shortcuts...!

OpenStudy (anonymous):

He's got one of them tutors....

OpenStudy (amistre64):

i wonder what is considered a shortcut :)

OpenStudy (amistre64):

is memorizing the formulas a shortcut?

OpenStudy (anonymous):

Makes u do the two sides separately..duh!

OpenStudy (anonymous):

@amistre64: sort of :)

OpenStudy (amistre64):

lol .... um, is posting it on openstudy a shortcut?

OpenStudy (anonymous):

i dont think so..?

OpenStudy (anonymous):

Designer shortcut...

OpenStudy (anonymous):

\[\tan \theta=\sin\theta/\cos\theta=(2\cos\theta\sin\theta)/(2\cos^{2}\theta)\](mulitply top and bottom by 2cos(theta) \[=\sin(2\theta)/[1+\cos^{2}\theta-(1-\cos^{2}\theta)]=\sin(2\theta)/(1+\cos^{2}\theta-\sin^{2}\theta)=\sin(2\theta)/[1+\cos(2\theta)]\]

OpenStudy (anonymous):

Gotta be worth a medal, that...

OpenStudy (anonymous):

lol...would've been easier to work right to left :)

OpenStudy (anonymous):

i know... but my teacher requires the left side... thank you though

OpenStudy (anonymous):

someone else can do #2 and #3 ;)

OpenStudy (anonymous):

okay.. thanks a lot! :D

OpenStudy (anonymous):

Like i said...on 2, go from right to left then reverse your steps...it's much easier to just apply the double angle than to build it. #3, go left to right usding the double angle formula dn it'll work out fine.

OpenStudy (anonymous):

in fact, i solved #1 going right to left and then typed my steps in backwards. :)

OpenStudy (anonymous):

why do you multiply 2cos theta to sin/cos?

OpenStudy (anonymous):

to build the cos squared on the bottom. i needed to get cos^2-sin^2 on the bottom to be able to use the double-angle

OpenStudy (anonymous):

that's the exact reason why it's often easier to work backwards...the motivation just isn't then when you have to "build" the formula instead of using it.

OpenStudy (anonymous):

ahhh... i'm not very good at using the "multiply by 1" method.. thanks for clarifying things out for me :D

OpenStudy (anonymous):

i think i can answer no 2 and 3 now.. thank you

OpenStudy (anonymous):

no problem...if you start working the others and get stuck, let us know. post as far as you get and someone here can nudge you on.